# Topology-Driven Parallel Trajectory Optimization in Dynamic Environments

### v1

@@ V-D Sensitivity Studies
V-D1 Sensitivity to the Number of Trajectories $P$
@@
V-D2 Sensitivity to the Consistency Parameter $c_{i}$
@@
TABLE III: Quantative results for simulations with uncertain obstacle motion of Sec. V-E over $200$ experiments. Task duration and runtime are reported as “mean (std. dev.)”. Best performance in bold.
# Method Dur. [s] Safe (%) Runtime [ms]
High Risk CC-MPC [4] 15.8 (2.3) 91 17.0 (8.7)
TCC-MPC++ (ours) 14.1 (0.8) 96 34.4 (7.9)
Medium Risk CC-MPC [4] 16.5 (2.7) 92 17.4 (9.6)
TCC-MPC++ (ours) 15.1 (1.4) 93 35.8 (8.2)
Low Risk CC-MPC [4] 17.2 (2.5) 90 18.4 (10.2)
TCC-MPC++ (ours) 16.1 (1.3) 97 38.1 (7.9)
V-E T-MPC Under Obstacle Uncertainty
@@
We deploy both planners in the scenario with $12$ randomized pedestrians and validate the planners under high ($\epsilon=0.1$), medium ($\epsilon=0.01$) and low ($\epsilon=0.001$) risk settings. The results are shown in Table III and visualized in Fig. 12. TCC-MPC++ consistently outperforms CC-MPC in isolation, leading to faster and more consistent task completion and fewer collisions. The local planner often collides when it becomes infeasible since it cannot recover. The initialization provided by the guidance planner resolves infeasibility and improves robustness.
@@ VI-A Experimental Setup
Figure 13: Overlayed top view camera images of real-world experiments of GT-MPC++ with one pedestrian. Blue and green overlays denote the robot and pedestrian trajectories, respectively. Timestamps of each image denoted in the upper left corner.
@@ VI-C Five Pedestrians
In the second set of experiments, we run LMPCC, T-MPC++, CC-MPC and TCC-MPC++ for $5$ minutes each with $5$ pedestrians in the space. Several runs of CC-MPC and TCC-MPC++ are visualized in Fig. 14 and Fig. 15, respectively. The reader is encouraged to watch the associated video [38]. Fig. 13(a) highlights a case where CC-MPC conservatively avoids two pedestrians, not detecting the pathway in between the pedestrians. In Fig. 13(b), CC-MPC gets infeasible and is not able to replan fast enough, requiring the robot to wait for a pedestrian to pass. In all experiments of Fig. 15, the robot passes pedestrians efficiently and smoothly. Fig. 14(a) and 14(b) highlight cases where the planner has to navigate through the crowd and does so successfully. In Fig. 14(c), we observe the planner falling in line with a pedestrian to pass another pedestrian.
The $5$ participants were not aware of the planner running in each experiment and the order of planners was randomized. After the experiments, the participants unanimously preferred the planners running in experiments $1$ and $4$, which were both guided planners. In general, participants reported that the guided planners felt safer and more predictable than the non-guided planners.
@@ VII-A Safety in Dynamic Environments
Planners deployed in the real world must be safe (i.e., collisions are unacceptable) and should not impede humans more than necessary. Table II indicated that all planners collided at least once in simulation. To identify the source of collisions, we repeated the experiments using the social forces model for both the pedestrian simulation model and the prediction model of the planner. The results are summarized in Table IV. Collisions are almost reduced to zero for T-MPC++ in this case, which shows that prediction mismatch causes most of the collisions.
TABLE IV: Quantative results for the simulations of Sec. V-C repeated with the pedestrian motion predictions following the social forces model.
# Method Task Duration [s] Safe (%) Runtime [ms]
@@
In the real-world experiments, we did not observe collisions. Pedestrians take direct evasive action when the robot impedes their safety, deviating from the social forces model. Collisions in simulation therefore seem to correspond to cases where the human must take evasive action in practice.
@@ VII-B Advantages of Parallel Optimization
VII-C Limitations and Future Work
One of the remaining limitations of the framework is the lack of interaction between the humans and the robot. It may be possible to reduce the complexity of interaction with humans to an explicit decision on the topology class of interaction (along the lines of [39]) that simplifies the planning problem. Future work could also explore if the consistency of the planner can be further improved by reacting to the appearance of new trajectories and the merger of existing trajectories. Finally, the guidance planner can possibly be extended to incorporate non-Gaussian (i.e., multi-modal) uncertainty in obstacle motion (e.g., [15]).
@@ VIII CONCLUSION
We simulated a mobile robot navigating among pedestrians and showed that the proposed guided planner resulted in faster, safer and more consistent robot motion than existing planners, including a state-of-the-art topology-guided planner. We qualitatively observed the same improvement in the real world, where we navigated successfully among five pedestrians.

### v2

@@ V-E Sensitivity Studies
V-E1 Sensitivity to the Number of Trajectories $P$
@@
V-E2 Sensitivity to the Consistency Parameter $c_{i}$
@@
V-F Empirical Cost Comparison
To verify that T-MPC++ is able to find lower-cost local optima than the local planner in isolation (as discussed in Sec. IV-F), we compare the optimal cost of the executed trajectory attained by LMPCC (local planner), T-MPC and T-MPC++ that use identical cost functions. We perform this study in the crowded environment of Sec. V-D over $50$ experiments. Table IV indicates that T-MPC finds lower-cost local optima than LMPCC and does so consistently (lower std. dev.). T-MPC++ further reduces this cost and its deviation. The average cost of T-MPC++ is less than half that of the non-guided planner.
TABLE IV: Comparison of the attained cost in the crowded environment of Sec. V-D over $50$ experiments. Notation, including the cost that is tested for significance, follows that of Table II. The cost excludes infeasible planner iterations.
# Ped. Method Dur. [s] Safe (%) Cost Runtime [ms]
50 LMPCC [3] 21.9 (1.7) 84 2.25 (9.36) 7.1 (4.8)
T-MPC (ours) 21.6 (1.5) 84 1.60 (6.21) 20.6 (7.7)
T-MPC++ (ours) 21.0 (0.9) 96 1.05 (4.67) 21.7 (7.4)
V-G T-MPC Under Obstacle Uncertainty
@@
We deploy both planners in the scenario with $12$ randomized pedestrians and compare the planners under high ($\epsilon=0.1$), medium ($\epsilon=0.01$) and low ($\epsilon=0.001$) risk settings. The results are shown in Table V and visualized in Fig. 11. TCC-MPC++ consistently outperforms CC-MPC in isolation, leading to significantly faster and more consistent task completion and fewer collisions. The local planner often collides when it becomes infeasible since it cannot recover. The initialization provided by the guidance planner resolves infeasibility and improves robustness.
TABLE V: Quantative results for simulations with uncertain obstacle motion of Sec. V-G over $200$ experiments. Notation follows that of Table II.
# Method Task Duration [s] Safe (%) Runtime [ms]
High Risk CC-MPC [4] 15.8 (2.3) 91 17.0 (8.7)
TCC-MPC++ (ours) 14.1 (0.8) 96 34.5 (7.9)
Medium Risk CC-MPC [4] 16.5 (2.7) 92 17.4 (9.7)
TCC-MPC++ (ours) 15.1 (1.4) 93 35.8 (8.2)
Low Risk CC-MPC [4] 17.2 (2.5) 90 18.5 (10.2)
TCC-MPC++ (ours) 16.1 (1.3) 97 38.1 (7.9)
Figure 11: Visualization of the task duration in Table V.
@@ VI-A Experimental Setup
Figure 12: Overlayed top view camera images of real-world experiments of TCC-MPC++ with one pedestrian. Blue and green overlays denote the robot and pedestrian trajectories, respectively. Timestamps of each image denoted in the upper left corner.
@@ VI-C Five Pedestrians
In the second set of experiments, we run LMPCC, T-MPC++, CC-MPC and TCC-MPC++ for $5$ minutes each with $5$ pedestrians in the space. The participants were instructed to walk naturally toward a point on the other side of the lab. Before starting the recorded experiments, participants were asked to walk for $3$ minutes without the robot, to get used to the environment. The participants were not aware of the planner running in each experiment and the order of planners was randomized. Several runs of CC-MPC and TCC-MPC++ are visualized in Fig. 13 and Fig. 14, respectively. The reader is encouraged to watch the associated video [45]. Fig. 13(a) highlights a case where CC-MPC conservatively avoids two pedestrians, not detecting the pathway in between the pedestrians. In Fig. 13(b), CC-MPC gets infeasible and is not able to replan fast enough, requiring the robot to wait for a pedestrian to pass. In all experiments of Fig. 14, the robot passes pedestrians efficiently and smoothly. Fig. 14(a) and 14(b) highlight cases where the planner has to navigate through the crowd and does so successfully. In Fig. 14(c), we observe the planner falling in line with a pedestrian to pass another pedestrian.
After the experiments, the participants unanimously preferred the planners running in experiments $1$ and $4$, which were both guided planners. In general, participants reported that the guided planners felt safer and more predictable than the non-guided planners.
@@ VII-A Safety in Dynamic Environments
Planners deployed in the real world must be safe (i.e., collisions are unacceptable) and should not impede humans more than necessary. Table II indicated that all planners collided at least once in simulation. To identify the source of collisions, we repeated the experiments using the social forces model for both the pedestrian simulation model and the prediction model of the planner (i.e., removing prediction mismatch). The results are summarized in Table VI. Collisions are almost reduced to zero for T-MPC++ in this case, which shows that prediction mismatch causes most of the collisions.
TABLE VI: Quantative results for the simulations of Sec. V-C repeated with the pedestrian motion predictions following the social forces model.
# Peds. Method Task Duration [s] Safe (%) Runtime [ms]
@@
In the real-world experiments, we did not observe collisions. We analyzed the video of the experiments and visually annotated the following instances where pedestrians had to take evasive action: $3$ out of $63$ interactions using LMPCC, $4$ out of $61$ interactions using CC-MPC, 00 out of $61$ interactions for T-MPC++ and 00 out of $60$ interactions for TCC-MPC++. This indicates that pedestrians take direct evasive action when the robot impedes their safety, deviating from the social forces model. Collisions in simulation therefore seem to correspond to cases where the human must take evasive action in practice.
@@ VII-B Advantages of Parallel Optimization
VII-C Selection of the Homotopy Class
The decision-making in Sec. IV-E used the optimal costs of the local planners to decide which trajectory to execute and preferred the homotopy class of the last followed trajectory. We observed in practice that the robot stayed closer to the reference path and velocity, and passed pedestrians behind rather than in front when necessary. We additionally observed that due to measurement and prediction noise, making a more consistent decision led to better navigation. With high consistency, the robot switches behavior only if the new one is significantly better than the current one. In this way we can ensure a more robust estimation of the cost of the trajectory. While our proposed decision-making led to fast navigation, it ignored social norms. The decision-making could be made more socially compliant by learning to pick the homotopy class that humans take from data (see e.g., [46, 34, 47]). Finally, as noted by [8], homotopy classes merge and split when obstacles are passed or appear in the planning horizon. Reacting to these events could make the planner more responsive in practice.
VII-D Limitations and Future Work
One of the remaining limitations of the framework is the lack of interaction between the humans and the robot. The social forces model that we used in simulation is interactive but does not accurately model human-robot interactions. It may be possible to reduce the complexity of interaction with humans to an explicit decision on the topology class of interaction (along the lines of [48]) that simplifies the planning problem. Additionally, T-MPC++ can possibly be extended to $3$D navigation in dynamic environments, for instance using higher-dimensional H-signatures, and the guidance planner could be extended to incorporate non-Gaussian (i.e., multi-modal) uncertainty in obstacle motion (e.g., [15]).
@@ VIII CONCLUSION
We simulated a mobile robot navigating among pedestrians and showed that the proposed guided planner resulted in faster and more consistent robot motion than existing planners, including a state-of-the-art topology-guided planner. We qualitatively observed the same improvement in the real world, where we navigated successfully among five pedestrians.
@@
Appendix A HOMOTOPY COMPARISON
This appendix details and compares three implementations of homotopy comparison function (3) for $2$D motion planning in dynamic environments.
A-A H-signature
The H-signature [30] approximates homotopy classes by homology classes, formally defined as follows.
Definition 4.
Figure 15: Illustrating example for the H-signature [30]. The two trajectories $\bm{\tau}_{1}$ (blue) and $\bm{\tau}_{2}$ (green) are in distinct homology classes as the loop that they form contains obstacle skeleton (orange). In practice, link (3) is placed far away.
If two trajectories are homotopic, they are homologous. The reverse does not hold. To compute the H-signature in the $3$D space composed of $x,y$-position and time, each obstacle and its prediction is virtually modeled as a current-carrying wire (see Fig. 15). The H-signature of Obstacle $j$, $h^{j}(\bm{\tau})$ is defined by the virtual magnetic field resulting from Obstacle $j$’s current loop, integrated over trajectory $\bm{\tau}$. If two trajectories $\bm{\tau}_{1},\bm{\tau}_{2}$ (see Fig. 15) enclose the loop of obstacle $j$, then $h^{j}(\bm{\tau}_{1})\neq h^{j}(\bm{\tau}_{2})$. Hence, two trajectories are equivalent if $h^{j}(\bm{\tau}_{1})=h^{j}(\bm{\tau}_{2})\ \forall j$ and are distinct otherwise. We compute the H-signature for a finite-time state space, by constructing a $1$D skeleton of the obstacle prediction that is looped outside of the workspace and time horizon (see Fig. 15). The skeleton is composed of the following lines. 1 The obstacle’s prediction for $0<t<T$. 2 A line upwards to $t=T+\epsilon$ where $\epsilon>0$ is a small constant and a line going outside of the workspace. 3 A line down to $t=-\epsilon$. 4 A line to the obstacle position at $t=0$. This skeleton ensures that trajectories can only enclose the predicted obstacle motion for $0<t<T$.
We assume that obstacle trajectories are piecewise linear (e.g., discrete-time trajectories). The integration of the magnetic field $\bm{B}$ can then be computed analytically (see [30]) per segment $i$ of Obstacle $j$’s skeleton $\overline{\bm{o}_{i}^{j}\bm{o}_{i}^{j^{\prime}}}$ as follows:
$\bm{p}=\bm{o}_{i}^{j}-\bm{r},\ \bm{p}^{\prime}=\bm{o}_{i}^{j^{\prime}}-\bm{r},\ \bm{d}=\frac{(\bm{o}_{i}^{j^{\prime}}-\bm{o}_{i}^{j})\times(\bm{p}\times\bm{p}^{\prime})}{||\bm{o}_{i}^{j^{\prime}}-\bm{o}_{i}^{j}||^{2}},$
$\bm{\Phi}(\bm{o}_{i}^{j},\bm{o}_{i}^{j^{\prime}},\bm{r})=\frac{1}{||\bm{d}||^{2}}\left(\frac{\bm{d}\times\bm{p}^{\prime}}{||\bm{p}^{\prime}||}-\frac{\bm{d}\times\bm{p}}{||\bm{p}||}\right),$
$\bm{B}(\bm{r})=\frac{1}{4\pi}\sum_{i=0}^{I}\bm{\Phi}(\bm{o}_{i}^{j},\bm{o}_{i}^{j^{\prime}},\bm{r}),$
[…]
A-B Winding Number
The winding number [39] is a topological invariant that indicates how the robot and obstacle $j$ are rotated around each other. It is computed as follows. The relative position of obstacle $j$ to the robot for time step $k$ is $\bm{d}_{k}^{j}=\bm{p}_{k}-\bm{o}_{k}^{j}$. The relative angle $\angle\theta_{k}^{j}$ is the angle of $\bm{d}_{k}^{j}$ in a fixed global frame. Between time steps $k$ and $k+1$, the relative angle changes by $\Delta\theta_{k}^{j}=\theta_{k+1}^{j}-\theta_{k}^{j}$. The winding number accumulates these changes over all time steps, $\lambda(\bm{\tau},\bm{o}^{j})=\frac{1}{2\pi}\sum_{k=1}^{N}\Delta\theta_{k}^{j}.$ The sign of the winding number $\lambda$ indicates the passing direction, its magnitude denotes passing progress. We consider a trajectory to pass obstacle $j$ if $|\lambda^{j}|\geq\lambda_{\textrm{pass}}$, where by default $\lambda_{\textrm{pass}}=\frac{1}{4\pi}$. We consider two trajectories distinct if there exists at least one obstacle that the trajectories pass on different sides and consider them equivalent otherwise [^8] Future work could also use winding numbers to distinguish between passing and non-passing trajectories.. We cache computed winding numbers to prevent recomputation.
A-C Universal Visibility Deformation
Universal Visibility Deformation (UVD) was proposed for static obstacle avoidance in $3$D and therefore does not exactly capture the local optima for collision avoidance in $2$D dynamic environments. Two trajectories are in the same UVD class if points along the trajectories can be connected, without intersecting with obstacles.
Definition 5.
[33] Two trajectories $\bm{\tau}_{1}(s),\bm{\tau}_{2}(s)$ parameterized by $s\in[0,1]$ and satisfying $\bm{\tau}_{1}(0)=\bm{\tau}_{2}(0)$, $\bm{\tau}_{1}(1)=\bm{\tau}_{2}(1)$, belong to the same uniform visibility deformation class, if for all $s$, line $\overline{\bm{\tau}_{1}(s)\bm{\tau}_{2}(s)}$ is collision-free.
In practice, we check collisions for $s$ at discrete intervals along the trajectories.
TABLE VII: Comparison between homotopy comparison implementations. Notation follows that of Table II.
# Method Dur. [s] Safe (%) Homotopy Comparison Time [ms]
Homotopy Comparison Winding Angles 21.2 (0.9) 93 0.3 (0.7)
H-Signature 21.2 (1.0) 92 2.1 (4.0)
UVD 21.1 (0.9) 88 3.7 (8.5)
A-D Comparison
We compare the homotopy comparison functions in simulation on the scenario of Sec. V-D over $100$ experiments. Table VII indicates that UVD degrades navigation performance, likely because it is not designed for dynamic environments and may lead to duplicate trajectories in practice. The H-signature and winding numbers show similar navigation performance. Winding numbers are computationally more efficient but require a minimum passing angle to be tuned. Since the H-signature generalizes to higher dimensions and both methods are still real-time, we opted to use the H-signature in this paper.